OpenStudy (anonymous):

Solve for x : (x^2+2)^2 + 8x^2 = 6x(x^2+2) The answers given are 1 ± i , 2 ± √2

OpenStudy (anonymous):

\[\frac{-[-(3+\sqrt{17})]\pm \sqrt{(3+\sqrt{17})^2-4(1)(2)} }{ 2(1) }\]

OpenStudy (anonymous):

\[\frac{-[-(3+\sqrt{17})]\pm \sqrt{9+6\sqrt{17}+17-8} }{ 2 }\]

OpenStudy (anonymous):

\[\frac{3+\sqrt{17}\pm \sqrt{6(3+\sqrt{17})} }{ 2 }\]

OpenStudy (anonymous):

@ksananthu : is this what you meant ?

OpenStudy (anonymous):

i'm sorry there is a mistake in my calculation

OpenStudy (anonymous):

wait i will correct now

OpenStudy (anonymous):

(x^2 + 2)^2 - 2*3 x (x^2 + 2) + 8 x^2 = 0 {(x^2 + 2)^2 - 2*3 x (x^2 + 2) + 9 x^2} +8 x^2 - 9 x^2 = 0 {(x^2 + 2) - 3 x}^2 - x^2 = 0 {(x^2 + 2) - 3 x}^2=x^2 taking roots bothsides (x^2 + 2) - 3 x = x x^2 - 4 x+2=0 \[x=\frac{ 4\pm \sqrt{16-4*1*2} }{ 2 }=2\pm \sqrt{2}\]

OpenStudy (anonymous):

or by taking roots we get another equation also (x^2 + 2) - 3 x =- x x^2 - 2 x+2=0 then it will give the value of x=1±i

OpenStudy (anonymous):

now you got the answer

OpenStudy (anonymous):

thanks :)